A capacitor of 60 C is connected to a 200 V supply and another capacitor of 30 C is connected to 100V. Find the charge and energy stored on each capacitor. Now these two capacitors are disconnected from the supply and are joined to each other such that positive plate of first is connected to the positive plate of second and similarly the negative plates of both are joined together. Find the Loss in energy of the system. Why does this loss occurs?

Grade:12

1 Answers

kaushik dey
5 years ago
FOR THIS TYPE OF PROBLEM GUIDING PRINCIPLE IS : CHARGE IS CONSERVED
 
INITIAL CHARGE: 200X60+30X100= 15X1000
 
AFTER CHANGING CONFIGURATION BOTH CAPACITANCE IN PARALLEL
SO EQUIVALENT CAP= 90
 
LET COMMON TERMINAL VOLTAGE BE V
 
AS CHARGE CONSERVED
 
90XV= 15000
V= 15000/90= 500/3
 
ONCE THIS VOLTAGE IS FOUND
 
CHANGE IN ENERGY  CAN BE CALCUALTED BY USING ½ CV2  FORMULA. PLEASE DUE IT YOURSELF
 
LOSS OF ENERGY IS DUE TO ENEGRY LOSSES IN CHARGE RE-DISTRIBUTION BETWEEN CAPACITORS.